If a sprinter’s mass is [Number 53.4] kg, how much forward force must be exerted on the sprinter to make the sprinter accelerate at [Number 0.7] m/s2?
Answer #1

mass of the sprinter , m = 53.4 kg

acceleration of the sprinter , a = 0.7 m/s2

Forward force to be exerted on the sprinter, F = m*a = (53.4 * 0.7 ) = 37.38 N (ans)